# The discrete periodic Pitman transform: invariances, braid relations, and Burke properties

### v2

@@ 2.2. Relation to the periodic polymer
Let $\mathbf{W}=(\mathbf{W}_{i})_{i\in\mathbb{Z}}\in(\mathbb{Z}_{N})^{\mathbb{Z}}$ and $k,\ell\in\mathbb{Z}$ such that $\ell\geq k$, then:
$\Bigl{(}H(\mathbf{W}_{k})H(\mathbf{W}_{k+1})\ldots H(\mathbf{W}_{\ell})\Bigr{)}_{(i,j)}=Z^{\mathbf{W}}(\ell,j\mid k,i)\quad\forall i,j\in\mathbb{Z}.$
In other words, the partition function across any two points $(k,i)$ and $(\ell,j)$ on the directed lattice $\mathbb{Z}^{2}$ with periodic weights $\mathbf{W}$ is given by the $(i,j)$th entry in the matrix product of the $H$ matrices.
@@
$H(\mathbf{W}_{k})_{(i,j)}=\begin{cases}\exp\Biggl{(}{\sum\limits_{r=i}^{j}}W_{(k,r)}\Biggr{)},\quad&j\geq i\\ 0,\ &\text{otherwise}\end{cases}$
@@
$Z^{W}(k,j\mid k,i)=\begin{cases}\exp\Biggl{(}{\sum\limits_{r=i}^{j}}W_{(k,r)}\Biggr{)},\quad&j\geq i\\ 0,\ &\text{otherwise}\end{cases}=H(\mathbf{W}_{k})_{(i,j)}$
@@
$Z^{\mathbf{W}}(\ell+1,j\mid k,i)$ $=\sum\limits_{r=i}^{j}\Biggl{(}Z^{\mathbf{W}}(\ell,r\mid k,i)\exp\bigl{(}\sum_{s=r}^{j}W_{(\ell+1,s)}\bigr{)}\Biggr{)}$
$=\sum\limits_{r=i}^{j}\Biggl{(}Z^{\mathbf{W}}(\ell,r\mid k,i)\Bigl{(}H(\mathbf{W}_{\ell+1})\Bigr{)}_{(r,j)}\Biggr{)}.$
@@
$\sum\limits_{r=i}^{j}\Biggl{(}\Bigl{(}H(\mathbf{W}_{k})H(\mathbf{W}_{k+1})\ldots H(\mathbf{W}_{\ell})\Bigr{)}_{(i,r)}\Bigl{(}H(\mathbf{W}_{\ell+1})\Bigr{)}_{(r,j)}\Biggr{)}.$
Remembering now that $\bigl{(}H(\mathbf{W}_{\ell+1})\bigr{)}_{(r,j)}=0$ for $r>j$, and $\bigl{(}H(\mathbf{W}_{k})H(\mathbf{W}_{k+1})\ldots H(\mathbf{W}_{\ell})\bigr{)}_{(i,r)}=0$ for $r<i$, we see that
$\quad\;Z^{\mathbf{W}}(\ell+1,j\mid k,i)$
$=\sum\limits_{r=-\infty}^{\infty}\Biggl{(}\Bigl{(}H(\mathbf{W}_{k})H(\mathbf{W}_{k+1})\ldots H(\mathbf{W}_{\ell})\Bigr{)}_{(i,r)}\Bigl{(}H(\mathbf{W}_{\ell+1})\Bigr{)}_{(r,j)}\Biggr{)}$
$=\Bigl{(}H(\mathbf{W}_{k})H(\mathbf{W}_{k+1})\ldots H(\mathbf{W}_{\ell})H(\mathbf{W}_{\ell+1})\Bigr{)}_{(i,j)}.\qed$
We now show the invariance of polymer partition function under the periodic Pitman transform. Retaining the periodic weight vectors $\mathbf{W}=(\mathbf{W}_{r}\in\mathbb{R}^{\mathbb{Z}_{N}}:r\in\mathbb{Z})$ as before, for $k\in\mathbb{Z}$, recall the operator $\mathcal{P}_{k}:(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}\to(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}$ where
$\mathcal{P}_{k}\mathbf{W}=(\ldots,\mathbf{W}_{k-1},T(\mathbf{W}_{k+1},\mathbf{W}_{k}),D(\mathbf{W}_{k+1},\mathbf{W}_{k}),\mathbf{W}_{k+2,\ldots}).$
@@
$Z^{\mathcal{P}_{k_{m}}\cdots\mathcal{P}_{k_{1}}\mathbf{W}}(c,d\mid a,b)=Z^{\mathbf{W}}(c,d\mid a,b).$
@@
$Z^{\mathbf{W}}(c,d\mid a,b)=\bigl{(}H(\mathbf{W}_{a})H(\mathbf{W}_{a+1})\cdots H(\mathbf{W}_{c})\bigr{)}_{(b,d)},$
which we claim is preserved by the action $\mathcal{P}_{k_{m}}\ldots\mathcal{P}_{k_{1}}$ for any $b,d\in\mathbb{Z}$. In our base case $n=1$, we split into three sub-cases: $k_{1}<a-1$, $k_{1}>c$, and $a\leq k_{1}<c$. In the first two sub-cases, $\mathbf{W}_{a}\ldots\mathbf{W}_{c}$ are unaffected by $\mathcal{P}_{k_{1}}$, so our invariance holds trivially.
If $a\leq k_{1}<c$, then Corollary 2.5 gives that
$Z^{\mathbf{W}}(c,d\mid a,b)$ $=\bigl{(}H(\mathbf{W}_{a})H(\mathbf{W}_{a+1})\ldots H(\mathbf{W}_{k_{1}})H(\mathbf{W}_{k_{1}+1})\ldots H(\mathbf{W}_{c})\bigr{)}_{(b,d)}$
@@
$=Z^{\mathcal{P}_{k_{1}}\mathbf{W}}(c,d\mid a,b).$
Now, we assume by way of induction that the invariance holds for m1m-1 transforms. Again, if $k_{m}<a-1$ or $k_{m}>c$ our equivalence holds trivially. Otherwise, let $\mathbf{W}^{\prime}=\mathcal{P}_{k_{m-1}}\ldots\mathcal{P}_{k_{1}}\mathbf{W}$. Then
$Z^{\mathbf{W}}(c,d\mid a,b)$ $=Z^{\mathbf{W}^{\prime}}(c,d\mid a,b)$
$=Z^{\mathcal{P}_{k_{m}}\mathbf{W}^{\prime}}(c,d\mid a,b)$
$=Z^{\mathcal{P}_{k_{m}}\mathcal{P}_{k_{m-1}}\ldots\mathcal{P}_{k_{1}}\mathbf{W}}(c,d\mid a,b).$
@@
With Proposition 2.7 proved, we now prove Theorem 1.4 by extending to the multi-path result.
Proof of Theorem 1.4..
We use the Lindström-Gessel-Viennot (LGV) lemma (cf. [AZ18, Chapter 32], originally from [Lin73, GV85]) to prove the theorem. Because LGV applies only to edge-weighted graphs, we need to transfer our vertex weights to edge weights in the following way: for each vertex $(r,s)$, split it into $(r,s)_{\text{in}}$ and $(r,s)_{\text{out}}$, with edge $\bigl{(}(r,s)_{\text{in}},(r,s)_{\text{out}}\bigr{)}$ given the weight $\omega_{(r,s)}:=e^{W_{(r,s)}}$, and edges $\bigl{(}(r,s)_{\text{out}},(r,s+1)_{\text{in}}\bigr{)}$ and $\bigl{(}(r,s)_{\text{out}},(r+1,s)_{\text{in}}\bigr{)}$ given the weight 1. See Figure 5. Any path which started at $(r,s)$ starts at $(r,s)_{\text{in}}$, while any path ending at $(r,s)$ ends at $(r,s)_{\text{out}}$. Under this graph transformation, our partition function on the pre-transformed graph equals the sum over weight products along possible paths on the transformed diagram, so we can apply LGV.
[…]
Figure 5. Local transformation of a vertex-weighted node (left) with weight $W_{(r,s)}$ into an edge-weighted directed graph (right), where the internal edge carries weight $\omega_{(r,s)}=\exp(W_{(r,s)})$ and external edges carry weight 1. Paths in the original graph correspond to paths from “in” nodes to “out” nodes in the transformed graph, preserving path weights.
Write the path matrix $M=\bigl{(}Z^{\mathbf{W}}(c_{j},d_{j}\mid a_{i},b_{i})\bigr{)}_{(i,j)\in\llbracket 1,m\rrbracket}$. By LGV, we have
$\det M=\sum\limits_{P=(P_{1}\ldots P_{m})\in\Pi_{U,V_{\sigma}}\ \forall\sigma\in S_{m}}\operatorname{sgn}(\sigma)\prod\limits_{(i,j)\in P}e^{W_{(i,j)}}.$
By definition of $\Psi$, for $\sigma\neq\text{Id}$, $\Pi_{U,V_{\sigma}}=\varnothing$. Because $\operatorname{sgn}(\text{Id})=1$, we have that
$\det M=\sum\limits_{P=(P_{1}\ldots P_{m})\in\Pi_{U,V}}\prod\limits_{(i,j)\in P}e^{W_{(i,j)}}=Z^{\mathbf{W}}(V\mid U).$
By Proposition 2.7, each $M_{i,j}=Z^{\mathbf{W}}(c_{j},d_{j}\mid a_{i},b_{i})$ remains invariant under $\mathcal{P}_{k_{m}}\cdots\mathcal{P}_{k_{1}}$, so $M$ is unchanged, and so $\det M$ is invariant. Thus, $Z^{\mathbf{W}}(V\mid U)=\det M$ is invariant. ∎
3. Braid relations: proof of Theorem 1.6
In this section, we prove the braid relation stated in Theorem 1.6. Notice that this relation translates into the following three equalities:
$T\Bigl{(}T\bigl{(}\mathbf{W}_{k+2},D(\mathbf{W}_{k+1},\mathbf{W}_{k})\bigr{)},T(\mathbf{W}_{k+1},\mathbf{W}_{k})\biggr{)}=T\Bigl{(}T(\mathbf{W}_{k+2},\mathbf{W}_{k+1}),\mathbf{W}_{k}\Bigr{)},$ (3.1)
$D\Bigl{(}T\bigl{(}\mathbf{W}_{k+2},D(\mathbf{W}_{k+1},\mathbf{W}_{k})\bigr{)},T(\mathbf{W}_{k+1},\mathbf{W}_{k})\Bigr{)}=T\Bigl{(}D(\mathbf{W}_{k+2},\mathbf{W}_{k+1}),D\bigl{(}T(\mathbf{W}_{k+2},\mathbf{W}_{k+1}),\mathbf{W}_{k}\bigr{)}\Bigr{)},$ (3.2)
$D\Bigl{(}\mathbf{W}_{k+2},D(\mathbf{W}_{k+1},\mathbf{W}_{k})\Bigr{)}=D\Bigl{(}D(\mathbf{W}_{k+2},\mathbf{W}_{k+1}),D\bigl{(}T(\mathbf{W}_{k+2},\mathbf{W}_{k+1}),\mathbf{W}_{k}\bigr{)}\Bigr{)}.$ (3.3)
Equation (3.3) follows from [CGS24, Proposition 5.13], as explained in Lemma A.8 of the appendix. We now use the partition function invariance, Theorem 1.4, to demonstrate that (3.1) implies both (3.2) and (3.3). The main tool will be the following lemma:

### v3

@@ 2.2.2. Proof of Item (ii)
For $\mathbf{W}=(\mathbf{W}_{i})_{i\in\mathbb{Z}},\mathbf{W}^{\prime}=(\mathbf{W}^{\prime}_{i})_{i\in\mathbb{Z}}\in(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}$ satisfying $\mathfrak{s}(\mathbf{W}_{1})=\mathfrak{s}(\mathbf{W}^{\prime}_{1})$ we have that $H(\mathbf{W}_{1})H(\mathbf{W}_{2})=H(\mathbf{W}^{\prime}_{1})H(\mathbf{W}_{2}^{\prime})$ if and only if $\mathbf{W}_{1}=\mathbf{W}^{\prime}_{1}$ and $\mathbf{W}_{2}=\mathbf{W}_{2}^{\prime}$.
Before proving Lemma 2.8, we first use Lemmas 2.8 and A.2 to complete the proof of Theorem 1.1 (ii):
Proof of Theorem 1.1(ii)..
We apply Proposition 2.7 three times: first to the last two terms, then the first two, then the last two again:
$H\bigl(\mathbf{X}_{k}\bigr)H\bigl(\mathbf{X}_{k+1}\bigr)H\bigl(\mathbf{X}_{k+2}\bigr)=$
[…]
$\qquad H\Bigl(D\bigl(D\bigl(\mathbf{X}_{k},T(\mathbf{X}_{k+1},\mathbf{X}_{k+2})\bigr),D(\mathbf{X}_{k+1},\mathbf{X}_{k+2})\bigr)\Bigr).$ (2.14)
Alternatively, we apply Proposition 2.7 to the first two terms, then the last two terms, then again the first two terms to get
$H\bigl(\mathbf{X}_{k}\bigr)H\bigl(\mathbf{X}_{k+1}\bigr)H\bigl(\mathbf{X}_{k+2}\bigr)=$
[…]
$\qquad H\Bigl(D\bigl(T(\mathbf{X}_{k},\mathbf{X}_{k+1}),T\bigl(D(\mathbf{X}_{k},\mathbf{X}_{k+1}),\mathbf{X}_{k+2}\bigr)\big)\Bigr)H\Bigl(D\bigl(D(\mathbf{X}_{k},\mathbf{X}_{k+1}),\mathbf{X}_{k+2}\bigr)\Bigr).$ (2.15)
$H\Bigl(T\bigl(\mathbf{X}_{k},T(\mathbf{X}_{k+1},\mathbf{X}_{k+2})\bigr))\Bigr)H\Bigl(T\bigl(D\bigl(\mathbf{X}_{k},T(\mathbf{X}_{k+1},\mathbf{X}_{k+2})\bigr),D(\mathbf{X}_{k+1},\mathbf{X}_{k+2})\bigr)\Bigr)$
$=H\Bigl(T\bigl(T(\mathbf{X}_{k},\mathbf{X}_{k+1}),T\bigl(D(\mathbf{X}_{k},\mathbf{X}_{k+1}),\mathbf{X}_{k+2}\bigr)\big)\Bigr)H\Bigl(D\bigl(T(\mathbf{X}_{k},\mathbf{X}_{k+1}),T\bigl(D(\mathbf{X}_{k},\mathbf{X}_{k+1}),\mathbf{X}_{k+2}\bigr)\big)\Bigr).$
By repeated application of Lemma 2.1, we have that
$\mathfrak{s}(T(\mathbf{X}_{k},T(\mathbf{X}_{k+1},\mathbf{X}_{k+2})))=\mathfrak{s}(\mathbf{X}_{k+2})=\mathfrak{s}(T(T(\mathbf{X}_{k},\mathbf{X}_{k+1}),T(D(\mathbf{X}_{k},\mathbf{X}_{k+1}),\mathbf{X}_{k+2}))).$
Proof of Lemma 2.8..
Immediately, the backwards direction is clear. Then, it remains to show that, if $H(\mathbf{W}_{1})H(\mathbf{W}_{2})=H(\mathbf{W}_{1}^{\prime})H(\mathbf{W}_{2}^{\prime})$, then $\mathbf{W}_{1}=\mathbf{W}_{1}^{\prime}$ and $\mathbf{W}_{2}=\mathbf{W}_{2}^{\prime}$. Let us prove by backwards induction on $k=N,\ldots,1$ that $W_{1,k}=W_{1,k}^{\prime}$ and $W_{2,k}=W_{2,k}^{\prime}$.
We start with $k=N$ as our base case. From Lemma 3.1 below, we have that
$\left(H(\mathbf{W}_{1})H(\mathbf{W}_{2})\right)_{1,N}$ $=Z^{\mathbf{W}}(2,N\mid 1,1)$
$=e^{W_{2,N}}\left(Z^{\mathbf{W}}(1,N\mid 1,1)+Z^{\mathbf{W}}(2,N-1\mid 1,1)\right)$
$=e^{W_{2,N}}\Biggl(\prod_{i=1}^{N}e^{W_{1,i}}+\left(H(\mathbf{W}_{1})H(\mathbf{W}_{2})\right)_{1,N-1}\Biggr).$ (2.16)
Similarly, we conclude that
$\left(H(\mathbf{W}^{\prime}_{1})H(\mathbf{W}^{\prime}_{2})\right)_{1,N}$ $=Z^{\mathbf{W}^{\prime}}(2,N\mid 1,1)$
$=e^{W^{\prime}_{2,N}}\left(Z^{\mathbf{W}^{\prime}}(1,N\mid 1,1)+Z^{\mathbf{W}^{\prime}}(2,N-1\mid 1,1)\right)$
$=e^{W^{\prime}_{2,N}}\Biggl(\prod_{i=1}^{N}e^{W^{\prime}_{1,i}}+\left(H(\mathbf{W}^{\prime}_{1})H(\mathbf{W}^{\prime}_{2})\right)_{1,N-1}\Biggr).$ (2.17)
By assumption, we have $H(\mathbf{W}_{1})H(\mathbf{W}_{2})=H(\mathbf{W}^{\prime}_{1})H(\mathbf{W}_{2}^{\prime})$ and $\prod_{i=1}^{N}e^{W_{1,i}}=e^{\mathfrak{s}(W_{1})}=e^{\mathfrak{s}(W^{\prime}_{1})}=\prod_{i=1}^{N}e^{W^{\prime}_{1,i}}$. Equating (2.16) and (2.17) and canceling the nonzero expressions in parentheses yields $W_{2,N}=W^{\prime}_{2,N}$. Next, we see that $W^{\prime}_{1,N}=W_{1,N}$, since:
$(H(\mathbf{W}_{1})H(\mathbf{W}_{2}))_{(N,N)}=Z^{\mathbf{W}}(2,N\mid 1,N)=e^{W_{1,N}+W_{2,N}},$ (2.18)
and
$(H(\mathbf{W}^{\prime}_{1})H(\mathbf{W}_{2}^{\prime}))_{(N,N)}=Z^{\mathbf{W}^{\prime}}(2,N\mid 1,N)=e^{W^{\prime}_{1,N}+W^{\prime}_{2,N}}.$ (2.19)
To finish the proof, simply notice that the proof of the above only required $\prod_{i=1}^{N}e^{W^{\prime}_{1,i}}=\prod_{i=1}^{N}e^{W_{1,i}}$. Indeed, if we have $\prod_{i=1}^{k}e^{W_{1,i}}=\prod_{i=1}^{k}e^{W_{2,i}}$ for some $k\in\llbracket 1,N\rrbracket$, then the same proof shows that $W_{1,k}=W_{1,k}^{\prime}$ and $W_{2,k}=W_{2,k}^{\prime}$. This follows by backwards induction on $k$, since
$\prod_{i=1}^{k}e^{W_{1,i}}=e^{\mathfrak{s}(\mathbf{W}_{1})}\Biggl(\prod_{i=k+1}^{N}e^{W_{1,i}}\Biggr)^{-1}.\qed$
3. Invariance of polymer partition functions
In this section, we prove Theorem 1.4. The key to our proof is a matrix encoding of polymer partition functions, analogous to that developed in [NY04]. In our setting, this is given by the following lemma.
Lemma 3.1.
Let $\mathbf{X}=(\mathbf{X}_{i})_{i\in\mathbb{Z}}\in(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}$ and $k,\ell\in\mathbb{Z}$ such that $\ell\geq k$, then:
$\Bigl(H(\mathbf{X}_{k})H(\mathbf{X}_{k+1})\cdots H(\mathbf{X}_{\ell})\Bigr)_{(i,j)}=Z^{\mathbf{X}}(\ell,j\mid k,i)\quad\forall i,j\in\mathbb{Z}.$
In other words, the partition function across any two points $(k,i)$ and $(\ell,j)$ on the directed lattice $\mathbb{Z}^{2}$ with periodic weights $\mathbf{X}$ is given by the $(i,j)$th entry in the matrix product of the $H$ matrices.
@@ 3. Invariance of polymer partition functions
$H(\mathbf{X}_{k})_{(i,j)}=\begin{cases}\exp\Biggl({\sum\limits_{r=i}^{j}}X_{(k,r)}\Biggr),\quad&j\geq i\\ 0,\ &\text{otherwise}\end{cases}$
@@
$Z^{X}(k,j\mid k,i)=\begin{cases}\exp\Biggl({\sum\limits_{r=i}^{j}}X_{(k,r)}\Biggr),\quad&j\geq i\\ 0,\ &\text{otherwise}\end{cases}=H(\mathbf{X}_{k})_{(i,j)}$
@@
$Z^{\mathbf{X}}(\ell+1,j\mid k,i)$ $=\sum\limits_{r=i}^{j}\Biggl(Z^{\mathbf{X}}(\ell,r\mid k,i)\exp\Bigl(\sum_{s=r}^{j}X_{(\ell+1,s)}\Bigr)\Biggr)$
$=\sum\limits_{r=i}^{j}\Biggl(Z^{\mathbf{X}}(\ell,r\mid k,i)\Bigl(H(\mathbf{X}_{\ell+1})\Bigr)_{(r,j)}\Biggr).$
@@
$\sum\limits_{r=i}^{j}\Biggl(\Bigl(H(\mathbf{X}_{k})H(\mathbf{X}_{k+1})\ldots H(\mathbf{X}_{\ell})\Bigr)_{(i,r)}\Bigl(H(\mathbf{X}_{\ell+1})\Bigr)_{(r,j)}\Biggr).$
Remembering now that $\bigl(H(\mathbf{X}_{\ell+1})\bigr)_{(r,j)}=0$ for $r>j$, and $\bigl(H(\mathbf{X}_{k})H(\mathbf{X}_{k+1})\ldots H(\mathbf{X}_{\ell})\bigr)_{(i,r)}=0$ for $r<i$, we see that
$Z^{\mathbf{X}}(\ell+1,j\mid k,i)$ $=\sum\limits_{r=-\infty}^{\infty}\Biggl(\Bigl(H(\mathbf{X}_{k})H(\mathbf{X}_{k+1})\ldots H(\mathbf{X}_{\ell})\Bigr)_{(i,r)}\Bigl(H(\mathbf{X}_{\ell+1})\Bigr)_{(r,j)}\Biggr)$
$=\Bigl(H(\mathbf{X}_{k})H(\mathbf{X}_{k+1})\ldots H(\mathbf{X}_{\ell})H(\mathbf{X}_{\ell+1})\Bigr)_{(i,j)}.\qed$
We now show the invariance of polymer partition function under the periodic Pitman transform. Retaining the periodic weight vectors $\mathbf{X}=(\mathbf{X}_{r}\in\mathbb{R}^{\mathbb{Z}_{N}}:r\in\mathbb{Z})$ as before, for $k\in\mathbb{Z}$, recall the operator $\mathcal{P}_{k}:(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}\to(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}$ where
$\mathcal{P}_{k}\mathbf{X}=(\ldots,\mathbf{X}_{k-1},T(\mathbf{X}_{k},\mathbf{X}_{k+1}),D(\mathbf{X}_{k},\mathbf{X}_{k+1}),\mathbf{X}_{k+2,\ldots}).$
@@
$Z^{\mathcal{P}_{k_{m}}\cdots\mathcal{P}_{k_{1}}\mathbf{X}}(c,d\mid a,b)=Z^{\mathbf{X}}(c,d\mid a,b).$
@@
$Z^{\mathbf{X}}(c,d\mid a,b)=\bigl(H(\mathbf{X}_{a})H(\mathbf{X}_{a+1})\cdots H(\mathbf{X}_{c})\bigr)_{(b,d)},$
which we claim is preserved by the action $\mathcal{P}_{k_{m}}\ldots\mathcal{P}_{k_{1}}$ for any $b,d\in\mathbb{Z}$. In our base case $m=1$, we split into three sub-cases: $k_{1}<a-1$, $k_{1}>c$, and $a\leq k_{1}<c$. In the first two sub-cases, $\mathbf{X}_{a}\ldots\mathbf{X}_{c}$ are unaffected by $\mathcal{P}_{k_{1}}$, so our invariance holds trivially.
If $a\leq k_{1}<c$, then Proposition 2.7 gives that
$Z^{\mathbf{X}}(c,d\mid a,b)$ $=\bigl(H(\mathbf{X}_{a})H(\mathbf{X}_{a+1})\ldots H(\mathbf{X}_{k_{1}})H(\mathbf{X}_{k_{1}+1})\ldots H(\mathbf{X}_{c})\bigr)_{(b,d)}$
@@
$=Z^{\mathcal{P}_{k_{1}}\mathbf{X}}(c,d\mid a,b).$
Now, we assume by way of induction that the invariance holds for m1m-1 transforms. Again, if $k_{m}<a-1$ or $k_{m}>c$ our equivalence holds trivially. Otherwise, let $\mathbf{X}^{\prime}=\mathcal{P}_{k_{m-1}}\ldots\mathcal{P}_{k_{1}}\mathbf{X}$. Then
$Z^{\mathbf{X}}(c,d\mid a,b)$ $=Z^{\mathbf{X}^{\prime}}(c,d\mid a,b)$
$=Z^{\mathcal{P}_{k_{m}}\mathbf{X}^{\prime}}(c,d\mid a,b)$
$=Z^{\mathcal{P}_{k_{m}}\mathcal{P}_{k_{m-1}}\ldots\mathcal{P}_{k_{1}}\mathbf{X}}(c,d\mid a,b).$
@@
With Proposition 3.2 proved, we now give one more intermediate lemma before proving Theorem 1.4.